# Virtual Multiplex Staining for Histological Images using a Marker-wise Conditioned Diffusion Model

### v1

@@ 4.1 Setup
We utilized two public datasets, HEMIT (Bian et al. 2024) and Orion-CRC (Lin et al. 2023), each providing paired H&E and multiplex-stained images, thus offering a robust benchmark for evaluating virtual multiplex staining techniques. (1) HEMIT contains H&E images and three mIHC markers: DAPI (highlighting cell nuclei), pan-cytokeratin (panCK, identifying tumor regions), and CD3 (marking T cells), all crucial for analyzing the tumor microenvironment and immune activity. The dataset consists of 5,292 $\times1024$ pixel image pairs with 50% overlap, split into training (3,717 pairs), validation (630 pairs), and test (945 pairs) sets. For rigorous evaluation, we curated 292 image pairs from the test set by excluding spatially overlapping regions and filtering out empty patches that can hinder meaningful quantitative evaluation. This strategy prevents artificial inflation of metrics (e.g., PSNR values above 60 dB) observed in biologically irrelevant, low-signal regions.
@@
pix2pix (Isola et al. 2017) 0.750 0.879 0.572 0.734 0.901 0.381 0.586 0.623 31.53 24.82 26.29 27.55
pix2pixHD (Wang et al. 2018) 0.702 0.844 0.579 0.709 0.948 0.592 0.726 0.755 33.36 26.14 28.07 29.19
HEMIT (Bian et al. 2024) 0.769 0.882 0.659 0.770 0.964 0.567 0.707 0.746 32.82 25.90 27.63 28.78
Parmar et al. (2024) 0.606 0.845 0.652 0.701 0.924 0.332 0.595 0.617 32.15 24.09 26.07 27.44
Marigold (Ke et al. 2024) 0.801 0.711 0.546 0.686 0.956 0.581 0.711 0.750 33.91 26.04 28.13 29.36
@@
We use the Structural Similarity Index Measure (SSIM), Pearson correlation score (R), and PSNR, following previous works (Burlingame et al. 2020; Bian et al. 2024).
@@
We use SD v2 as the backbone (Rombach et al. 2022), with patch sizes of $\times512$. Our models were trained and fine-tuned on four NVIDIA H100 GPUs following the settings of Ke et al. (2024) and Garcia et al. (2024). Further details are provided in the supplementary material.
@@
We compare our approach to several state-of-the-art methods, including pix2pix (Isola et al. 2017), pix2pixHD (Wang et al. 2018), HEMIT (Bian et al. 2024)Parmar et al. (2024), and Marigold (Ke et al. 2024). For the HEMIT dataset with 3 output channels, we implemented Marigold by injecting 3-channel mIHC images into the SD VAE encoder. pix2pixHD and Marigold were excluded from Orion-CRC evaluations due to architectural limitations for multi-channel (18-marker) generation. Pix2pix and HEMIT were adapted to output 18 marker channels for the Orion-CRC dataset. We retrained HEMIT on its own dataset after observing performance discrepancies with the authors’ released checkpoint. Parmar et al. was excluded from Orion-CRC evaluations due to severe mode collapse.
@@ 4.2 Results
On Orion-CRC, Table 2 demonstrates that our approach achieved the highest average scores, with improvements of +0.039 SSIM, +0.117 R, and +1.358 PSNR compared to the second-best results, achieving the highest SSIM for 13, R for 15, and PSNR for 18 out of 18 markers. For markers with sufficient signal presence, such as Hoechst (cellular structures), all methods performed well. HEMIT achieved the second-highest performance on Hoechst, but generally underperformed compared to pix2pix on other marker types, with pix2pix consistently showing the second-best performance overall. Our method showed lower R scores for CD31 and FOXP3, likely a consequence of the limited number of available patches for these markers (<<1,000). Nevertheless, as shown in Fig. 4, our model does not completely miss these signals and still provides reasonable localization (CD31 in row 2, FOXP3 in row 4); in contrast, other methods frequently miss signals for certain marker types such as SMA (row 2) and CD20 (row 5). While our results exhibit some mild blurriness and tend to predict certain markers over broader regions compared to the ground truth (e.g., panCK in row 2, CD68 and CD163 in row 3), this tendency is likely due to structural properties of the latent diffusion model, in particular the absence of skip connections and its operation in the latent space, as previously observed in (Parmar et al. 2024). For the remaining marker types, our method generally matches or exceeds the performance of previous approaches, yielding consistent improvements in both localization accuracy and overall metric scores.
@@
(DDIM steps: 50, Ensemble: 10 (Ke et al. 2024))
@@
Table 4 summarizes the impact of the fine-tuning stage on model performance and inference cost on the HEMIT dataset. A single-plex model (one model per marker) serves as baseline but is impractical for virtual multiplex staining with a large number of markers, requiring multiple models. Our multiplex model generates all markers in a single network, improving training efficiency and practical scalability. We used 10-fold test-time ensembling (Ke et al. 2024) for optimal performance. Multiplex model shows slightly higher SSIM and R, but with much heavier memory demand. Notably, single-step inference in the multiplex model yields even higher SSIM (0.757) than 50-step sampling with ensembling (0.673). After the fine-tuning stage, our approach achieves the highest performance with dramatically reduced runtime and memory, demonstrating that fine-tuning is crucial for both image fidelity and efficient, scalable virtual multiplex staining.
@@ Discussion.
While our framework achieves superior performance for virtual multiplex staining, several important considerations for future research remain. Although single-step inference significantly reduces per-sample cost, the computational cost for inference increases with the number of markers, which is an inherent challenge for high-dimensional multiplex applications. Moreover, our model does not explicitly utilize inter-marker correlations, potentially limiting multiplex image fidelity. Addressing these points will be important for realizing practical, large-scale virtual staining applications.
@@ 5 Conclusion
In this paper, we proposed a novel framework for virtual multiplex staining of histological images using a marker-wise conditioned diffusion model. By conditioning the diffusion process on marker-specific embeddings and leveraging pretrained LDM parameters, we enabled the generation of high-quality multiplex images from H&E-stained tissue. To further improve color fidelity, we refined our model with pixel-level loss functions, while reducing inference cost and improving practical applicability. Our method, validated on two public datasets, demonstrate state-of-the-art performance in generating multiplex marker images from a H&E section. In the future, we aim to enhance structure detail, explicitly model inter-marker correlations, and evaluate clinical applicability in pathology.
@@ 6 More Details on Datasets
Figure 7: Visualization of H&E tissue image patches from the Orion-CRC dataset and their corresponding binary cell segmentation masks (Lin et al. 2023). Each patch visualized is $\times256$ pixels in size. The segmentation masks were used to filter out patches with low cell content for experimental analysis.
@@ 6.2 Setup on Orion-CRC
Figure 7 shows example H&E tissue images from the Orion-CRC (Lin et al. 2023) dataset alongside their paired binary cell segmentation masks, each visualized at a size of $\times256$ pixels. During dataset preparation, these binary masks were used to filter patches. As a result, only patches with at least 25% of the area covered by cells were included in the experiment.
@@
Table 6: Number of patches per IF marker in the Orion-CRC (Lin et al. 2023) dataset used for quantitative analysis.
Figure 8: Comparison of pixel intensity distributions for different markers in an example image from the HEMIT (Bian et al. 2024) dataset. (a) Density plots of normalized pixel intensities, showing the distribution of pixel values in the range (0, 1] and the corresponding zero-ratios for each mIHC marker channel. (b) H&E image, (c) panCK (zero-ratio: 91%), (d) CD3 (zero-ratio: 87%), and (e) DAPI (zero-ratio: 0%).
@@ 6.3 Marker Type Distributions
Fig. 8 shows the pixel intensity characteristics of different markers in an example image from the HEMIT dataset (Bian et al. 2024). Density plot (a) quantitatively illustrates the percentage distribution of normalized pixel values (range: 0-1, excluding 0) and the proportion of background pixels (zero-ratio) for each marker channel of the example image. The example consists of an H&E image (b) paired with the corresponding mIHC marker images (c-e). The analysis reveals significant differences in marker expression patterns and background prevalence across markers. For clarity of visualization, the contrast and brightness of mIHC images (c-e) have been adjusted.
@@ 7.1 Marker-wise Conditional Diffusion Model
We followed the implementation settings of Ke et al. (2024) with an effective batch size of 32. The Adam optimizer was employed with an initial learning rate of 3e-5, a 100-step warm-up period, and exponential decay over 30,000 iterations. All experiments used four NVIDIA H100 GPUs. The DDPM noise scheduler (Ho, Jain, and Abbeel 2020) was set with $T=1000$ timesteps and incorporated multi-resolution noise (Whitaker 2024). Text conditioning was disabled by injecting an empty token.
@@ 7.2 Fine-tuning for Color Contrast Fidelity
The fine-tuning stage used the same learning rate schedule as above, starting at an initial learning rate of 3e-5 with a 100-step warm-up and exponential decay. Due to increased memory requirements for pixel-level loss backpropagation, batch sizes were reduced compared to 7.1.

### v2

@@ 4.1 Setup
We utilized two public datasets, HEMIT (bian2024hemit) and Orion-CRC (lin2023high), each providing paired H&E and multiplex-stained images, thus offering a robust benchmark for evaluating virtual multiplex staining techniques. (1) HEMIT contains H&E images and three mIHC markers: DAPI (highlighting cell nuclei), pan-cytokeratin (panCK, identifying tumor regions), and CD3 (marking T cells), all crucial for analyzing the tumor microenvironment and immune activity. The dataset consists of 5,292 $\times1024$ pixel image pairs with 50% overlap, split into training (3,717 pairs), validation (630 pairs), and test (945 pairs) sets. For rigorous evaluation, we curated 292 image pairs from the test set by excluding spatially overlapping regions and filtering out empty patches that can hinder meaningful quantitative evaluation. This strategy prevents artificial inflation of metrics (e.g., PSNR values above 60 dB) observed in biologically irrelevant, low-signal regions.
@@
pix2pix (isola2017image) 0.750 0.879 0.572 0.734 0.901 0.381 0.586 0.623 31.53 24.82 26.29 27.55
pix2pixHD (wang2018high) 0.702 0.844 0.579 0.709 0.948 0.592 0.726 0.755 33.36 26.14 28.07 29.19
HEMIT (bian2024hemit) 0.769 0.882 0.659 0.770 0.964 0.567 0.707 0.746 32.82 25.90 27.63 28.78
parmar2024one 0.606 0.845 0.652 0.701 0.924 0.332 0.595 0.617 32.15 24.09 26.07 27.44
Marigold (ke2024repurposing) 0.801 0.711 0.546 0.686 0.956 0.581 0.711 0.750 33.91 26.04 28.13 29.36
@@
We use the Structural Similarity Index Measure (SSIM), Pearson correlation score (R), and PSNR, following previous works (burlingame2020shift; bian2024hemit).
@@
We use SD v2 as the backbone (rombach2022high), with patch sizes of $\times512$. Our models were trained and fine-tuned on four NVIDIA H100 GPUs following the settings of ke2024repurposing and garcia2024fine. Further details are provided in the supplementary material.
@@
We compare our approach to several state-of-the-art methods, including pix2pix (isola2017image), pix2pixHD (wang2018high), HEMIT (bian2024hemit)parmar2024one, and Marigold (ke2024repurposing). For the HEMIT dataset with 3 output channels, we implemented Marigold by injecting 3-channel mIHC images into the SD VAE encoder. pix2pixHD and Marigold were excluded from Orion-CRC evaluations due to architectural limitations for multi-channel (18-marker) generation. Pix2pix and HEMIT were adapted to output 18 marker channels for the Orion-CRC dataset. We retrained HEMIT on its own dataset after observing performance discrepancies with the authors’ released checkpoint. Parmar et al. was excluded from Orion-CRC evaluations due to severe mode collapse.
@@ 4.2 Results
On Orion-CRC, Table 2 demonstrates that our approach achieved the highest average scores, with improvements of +0.039 SSIM, +0.117 R, and +1.358 PSNR compared to the second-best results, achieving the highest SSIM for 13, R for 15, and PSNR for 18 out of 18 markers. For markers with sufficient signal presence, such as Hoechst (cellular structures), all methods performed well. HEMIT achieved the second-highest performance on Hoechst, but generally underperformed compared to pix2pix on other marker types, with pix2pix consistently showing the second-best performance overall. Our method showed lower R scores for CD31 and FOXP3, likely a consequence of the limited number of available patches for these markers (<<1,000). Nevertheless, as shown in Fig. 4, our model does not completely miss these signals and still provides reasonable localization (CD31 in row 2, FOXP3 in row 4); in contrast, other methods frequently miss signals for certain marker types such as SMA (row 2) and CD20 (row 5). While our results exhibit some mild blurriness and tend to predict certain markers over broader regions compared to the ground truth (e.g., panCK in row 2, CD68 and CD163 in row 3), this tendency is likely due to structural properties of the latent diffusion model, in particular the absence of skip connections and its operation in the latent space, as previously observed in (parmar2024one). For the remaining marker types, our method generally matches or exceeds the performance of previous approaches, yielding consistent improvements in both localization accuracy and overall metric scores.
@@
(DDIM steps: 50, Ensemble: 10 (ke2024repurposing))
@@
Table 4 summarizes the impact of the fine-tuning stage on model performance and inference cost on the HEMIT dataset. A single-plex model (one model per marker) serves as baseline but is impractical for virtual multiplex staining with a large number of markers, requiring multiple models. Our multiplex model generates all markers in a single network, improving training efficiency and practical scalability. We used 10-fold test-time ensembling (ke2024repurposing) for optimal performance. Multiplex model shows slightly higher SSIM and R, but with much heavier memory demand. Notably, single-step inference in the multiplex model yields even higher SSIM (0.757) than 50-step sampling with ensembling (0.673). After the fine-tuning stage, our approach achieves the highest performance with dramatically reduced runtime and memory, demonstrating that fine-tuning is crucial for both image fidelity and efficient, scalable virtual multiplex staining.
@@ Discussion.
While our framework achieves superior performance for virtual multiplex staining, several challenges remain. Single-step inference reduces per-sample cost, but computational cost scales with marker count, which is an inherent challenge for high-dimensional applications. Our model also lacks explicit inter-marker correlation modeling, potentially limiting image fidelity. Addressing these points is crucial for practical deployment.
@@ 5 Conclusion
In this paper, we proposed a novel framework for virtual multiplex staining using marker-wise conditioned diffusion models. By conditioning on marker-specific embeddings and refining with pixel-level loss functions, we achieved state-of-the-art performance on public datasets while improving practical applicability. Future directions include enhancing structural detail, explicitly modeling inter-marker correlations, and evaluating clinical impact in pathology.
6 Acknowledgments
This work was supported in part by the National Research Foundation of Korea under Grant RS-2024-00349697 and Grant RS-2021-NR060143; in part by the Institute for Information and Communications Technology Planning and Evaluation under Grant IITP-2025-RS-2020-II201819; in part by the Technology Development Program funded by the Ministry of SMEs and Startups (MSS), South Korea, under Grant RS-2024-00437796; in part by the National Research Council of Science and Technology (NST) grant funded by Korean Government [Ministry of Science and Information and Communications Technology (MSIT)] under Grant GTL24031-000; and in part by Korea University Grant.
@@ 7 More Details on Datasets
Figure 7: Visualization of H&E tissue image patches from the Orion-CRC dataset and their corresponding binary cell segmentation masks (lin2023high). Each patch visualized is $\times256$ pixels in size. The segmentation masks were used to filter out patches with low cell content for experimental analysis.
@@ 7.2 Setup on Orion-CRC
Figure 7 shows example H&E tissue images from the Orion-CRC (lin2023high) dataset alongside their paired binary cell segmentation masks, each visualized at a size of $\times256$ pixels. During dataset preparation, these binary masks were used to filter patches. As a result, only patches with at least 25% of the area covered by cells were included in the experiment.
@@
Table 6: Number of patches per IF marker in the Orion-CRC (lin2023high) dataset used for quantitative analysis.
Figure 8: Comparison of pixel intensity distributions for different markers in an example image from the HEMIT (bian2024hemit) dataset. (a) Density plots of normalized pixel intensities, showing the distribution of pixel values in the range (0, 1] and the corresponding zero-ratios for each mIHC marker channel. (b) H&E image, (c) panCK (zero-ratio: 91%), (d) CD3 (zero-ratio: 87%), and (e) DAPI (zero-ratio: 0%).
@@ 7.3 Marker Type Distributions
Fig. 8 shows the pixel intensity characteristics of different markers in an example image from the HEMIT dataset (bian2024hemit). Density plot (a) quantitatively illustrates the percentage distribution of normalized pixel values (range: 0-1, excluding 0) and the proportion of background pixels (zero-ratio) for each marker channel of the example image. The example consists of an H&E image (b) paired with the corresponding mIHC marker images (c-e). The analysis reveals significant differences in marker expression patterns and background prevalence across markers. For clarity of visualization, the contrast and brightness of mIHC images (c-e) have been adjusted.
@@ 8.1 Marker-wise Conditional Diffusion Model
We followed the implementation settings of ke2024repurposing with an effective batch size of 32. The Adam optimizer was employed with an initial learning rate of 3e-5, a 100-step warm-up period, and exponential decay over 30,000 iterations. All experiments used four NVIDIA H100 GPUs. The DDPM noise scheduler (ho2020denoising) was set with $T=1000$ timesteps and incorporated multi-resolution noise (whitaker2024multi). Text conditioning was disabled by injecting an empty token.
@@ 8.2 Fine-tuning for Color Contrast Fidelity
The fine-tuning stage used the same learning rate schedule as above, starting at an initial learning rate of 3e-5 with a 100-step warm-up and exponential decay. Due to increased memory requirements for pixel-level loss backpropagation, batch sizes were reduced compared to 8.1.